OpenStudy (laurenmyers1098):

I don't understand how to graph this step function? Question is in the comments.

OpenStudy (laurenmyers1098):

\[f(x) = 2 \times \lceil x \rceil \] for the domain \[-2 < x \le 1\]

OpenStudy (jtvatsim):

Let me look up the notation again, I always get ceiling and floor confused. :)

OpenStudy (laurenmyers1098):

Okay, thanks :)

OpenStudy (jtvatsim):

My understanding is that this is the ceiling function, you can think of it as the "round-up" function.

OpenStudy (jtvatsim):

So, this means that if we have \[\lceil 0.4 \rceil = 1\] the result rounds to 1

OpenStudy (jtvatsim):

of course, an extreme case is when \[\lceil -1.4 \rceil = -1\] which is kind of weird since "up" is "down" in negative land... Think of "up" as "more positive". :) Make sense so far as to how the ceiling function works?

OpenStudy (laurenmyers1098):

Yeah, I do understand how the ceiling function works, but what confuses me is how to even begin graphing it. Graphing is definitely not my forte

OpenStudy (jtvatsim):

OK, that shouldn't be too bad.

OpenStudy (jtvatsim):

We only need to graph from -2 < x <=1.

OpenStudy (jtvatsim):

Let's start with the right side of the graph. When x = 1. Then we have 2*ceiling(1), which rounds to 2*1 = 2.

OpenStudy (jtvatsim):

The graph looks like this. 

OpenStudy (jtvatsim):

See? The x = 1, and the y = 2. Is that ok?

OpenStudy (jtvatsim):

Remember that f(x) creates the y value. :)

OpenStudy (jtvatsim):

We aren't done yet, there is more, but this is our first step.

OpenStudy (laurenmyers1098):

Oh okay, I think it's starting to make sense. So to make the graph am I just plugging in values that are between -2 and 1? Including 1 as a value because of the less than or equal to sign?

OpenStudy (jtvatsim):

Yes, that is the idea. You can plot an "open dot" for x = -2 just for a reference if it helps.

OpenStudy (jtvatsim):

Would you like to see the rest of the graph? Or do you want to try it on your own first and then check?

OpenStudy (laurenmyers1098):

I wanna try it and then have you check, if that's alright :)

OpenStudy (jtvatsim):

Fine with me. That's the best way to learn! Let me know when you have it. :)

OpenStudy (jtvatsim):

I'll attach what I got to this message, just so you have it.

OpenStudy (jtvatsim):

Message if you have any questions! Good work and good luck! :)

OpenStudy (laurenmyers1098):

@jtvatsim I ended up getting the same graph as yours! Thank you so much!

OpenStudy (jtvatsim):

Excellent good job, you earn a medal! :)